40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 31 17:30:01 2022, Yehonathan, Update, BHD, Green shutters fixed 
    Reply  Wed Jun 1 11:56:28 2022, yuta, Update, ALS, Shutter driver for GRY replaced Screenshot_2022-06-01_12-01-53.png
Message ID: 16884     Entry time: Wed Jun 1 11:56:28 2022     In reply to: 16883
Author: yuta 
Type: Update 
Category: ALS 
Subject: Shutter driver for GRY replaced 

[JC, Yuta]

We replaced a shutter driver for GRY since it stopped working this morning.
We replaced it with a free driver which was sitting on the ITMY table.
The reverse polarity issue of C1:AUX-GREEN_Y_Shutter was fixed by switching one of the switches of the driver from N.O. to N.C.

Also, "Toggle" button was added to IFO_ALIGN.adl so that we can toggle shutters easily to find TEM00. It runs /home/controls/Git/40m/scripts/ALS/ShutterToggler.py.
 

Quote:
 

The green Y shutter now works but in reverese, meaning that sending 1 to C1:AUX-GREEN_Y_Shutter closes the shutter and vice versa. This needs to be fixed.

 

Attachment 1: Screenshot_2022-06-01_12-01-53.png  37 kB  Uploaded Wed Jun 1 13:02:02 2022  | Hide | Hide all
Screenshot_2022-06-01_12-01-53.png
ELOG V3.1.3-